Grade 12th passMechanics

A spring constant 800 N/m has extension of 5 cm. The work done in stretching it from 5 cm to 10 cm is ________ J
  1. 16
  2. 8
  3. 32
  4. 24

Yogiraj
10 Years agoGrade 12th pass

1 Answer

PASUPULETI   GURU MAHESH
10 Years ago
given k=800n/m
  x2-x1=10-5=5cm=0.05m
 we know that  W=1/2.K( x2-x1)2
                             SIMPLIFY YOU GET THE ANSWER
 THANK  YOU